The product of three no in A.P is 224 and largest no is 7 times the smaller. Find the numbers
Answer by mananth(16946)

let the numbers be
a-d, a , a+d
a(a^2-d^2) =224
a+d = 7(a-d)
a+d = 7a-7d
6a=8d
a=8d/6 = 4d/3


a(a^2-d^2) =224
4d/3((4d/3)^2 -d^2)=224
d(16d^2/9 -d^2) =56*3
d(16d^2-9d^2)/9)=56*3
d(7d^2) = 56*3*9
7d^3= 56*3*9/7
d^3= 8*3*9
d=6
a=8
the numbers 2,8,14